OpenStudy (anonymous):

A bald eagle snatches a fish from a lake and flies to an altitude of 256 ft. The fish manages to squirm free and falls back down into the lake. Its height h in feet can be modeled by h(t) = 256 - 16t2, where t is the time in seconds. How many seconds will the fish fall before hitting the water?

OpenStudy (radar):

It will hit the water when h = 0 0=256 - 16t^2 16t^2=256 t^2=256/16 = 16 t = 4 seconds